A company that usually sells satellite TV equipment for $72 and two years of satellite TV service for $648 has a special, timCan you check my answer?
Answer #1

Percentage of equipment in total price

= 72/(72+648) = 72/720

= 10%

Sales revenue = 410 * 10%

= $41

Your answer is correct :) :) :)